McCormick & Company is considering establishing new products in a new factory in Largo, Maryland. The project is expected to last for eight years. To determine the right financing option, you need to determine the appropriate discount based on the weighted average cost of capital. The cost of equity is estimated using the capital asset pricing model. Cash flows are assumed to be steady, the nominal risk-free rate for the short-term US government treasury bills is 1.5 percent, the 10-year government bonds rate is 2.5 percent, and the inflation rate is 2.54 percent. What is the real risk-free rate? Then, assume a beta of 1.2 and a market return of 5 percent. What is the cost of equity?
Answer #1

Real risk free rate=1.025/1.0254-1=-0.039%

Cost of equity=risk free rate+beta*(market return-risk free rate)=2.5%+1.2*(5%-2.5%)=5.5%